OpenStudy (anonymous):

Lucy wants to find out the number of frogs in a pond near her house. She catches 54 of them, puts a leg band on each frog, and puts them back in the pond. The next week, she catches 70 frogs. Of those 70 frogs, 13 have a leg band. Based on this information, what is a reasonable estimate of the number of frogs in the pond?

OpenStudy (anonymous):

13/70 = 54/x x = total approximately